# Generative AI for Intent-Driven Network Management in 6G: A Case Study on Hierarchical Learning Approach

### v1

@@ Abstract
With the emergence of 6G, mobile networks are becoming increasingly heterogeneous and dynamic, necessitating advanced automation for efficient management. Intent-Driven Networks (IDNs) address this by translating high-level intents into optimization policies. Large Language Models (LLMs) can enhance this process by understanding complex human instructions to enable adaptive, intelligent automation. Given the rapid advancements in Generative AI (GenAI), a comprehensive survey of LLM-based IDN architectures in disaggregated Radio Access Network (RAN) environments is both timely and critical. This article provides such a survey, along with a case study on a hierarchical learning-enabled IDN architecture that integrates GenAI across three key stages: intent processing, intent validation, and intent execution. Unlike most existing approaches that apply GenAI in the form of LLMs for intent processing only, we propose a hierarchical framework that introduces GenAI across all three stages of IDN. To demonstrate the effectiveness of the proposed IDN management architecture, we present a case study based on the latest GenAI architecture named Mamba. The case study shows how the proposed GenAI-driven architecture enhances network performance through intelligent automation, surpassing the performance of the conventional IDN architectures.
Index Terms:
Intent processing, Intent validation, Intent execution, Generative AI, Hierarchical learning
@@ I Introduction
Sixth-Generation (6G) networks are anticipated to support a diverse set of user requirements and have more complex deployments [1]. Traditional network management methods, which are based on manual configurations and human expertise, will face challenges when optimizing and operating such complex networks. Furthermore, manual configurations are costly, prone to errors and not scalable [2]. Intent-Driven Network (IDN) management is emerging as a solution to these challenges [1]. An intent defines desired outcomes by specifying target metrics without detailing implementation steps to ensure both human comprehension and machine interpretability [4]. Advances in Natural Language Processing (NLP) help operators express these intents more effectively. This reduces manual work, lowers errors, and speeds up configuration changes and service deployments [4, 5].
Disaggregated Radio Access Networks (RANs), like Open RAN with its layered structure, are well-suited for IDN management. High-level controllers can interpret operator goals and check if the intent is achievable based on real-time conditions. Low-level controllers or functions then create the best network policies to meet these goals. This approach improves flexibility, scalability, and decision-making efficiency.
On the other hand, Generative Artificial Intelligence (GenAI) has led to the rapid advancement of Large Language Models (LLMs) and revolutionized the field of NLP. As they have an outstanding ability to understand complex human instructions in natural language, IDN management is leaning towards LLMs instead of using fixed Service Level Agreement (SLA)-based approaches. For example, Dzeparoska et al. have introduced a pipeline that utilizes an LLM to process intents into structured and policy-based abstractions, which are next linked with Application Programming Interfaces (APIs) for execution [6]. Also, in our previous work we have introduced an IDN management framework that uses LLM few-shot learning for intent processing from a human network operator [4]. While these works use LLMs for intent processing, it can be highly beneficial to introduce GenAI not only as an intent processor via LLM but also we can take IDN management to the next level by introducing GenAI solutions in the other two stages namely intent validation and intent execution.
Figure 1: Three-step methodology for intent-driven network management.
With the inspiration for introducing efficient GenAI-based IDN management schemes via hierarchical disaggregated RAN environments, this work first reviews the state-of-the-art approaches that use GenAI for IDN management. Next, we introduce a three-step methodology for IDN management, encompassing intent processing, validation, and execution (see Fig. 1 for details). After that, we present how these three key steps can be implemented using a hierarchical RAN architecture with the help of multiple GenAI algorithms. Lastly, we present a case study that shows the implementation of these steps. In particular, we process intents in a memory-efficient manner where an LLM is fine-tuned with a custom dataset using Quantized Low Rank Adaptation (QLoRA) [7]. For the intent validation, a transformer-based time series predictor has been used in the case study which predicts crucial network parameters in future time slots and based on the predictions validates intents. Lastly, for intent execution, we generate network policies using Mamba, a GenAI architecture based on state space models for its demonstrated effectiveness compared to the transformer architecture in terms of memory and computation efficiency [8].
The core contributions of this paper are as follows: it presents a precise, yet comprehensive survey of LLM-based IDN architectures, highlighting current limitations and identifying opportunities for integrating broader GenAI techniques across the stages of intent processing, validation, and execution. Furthermore, it introduces a three-step methodology for IDN management that encompasses these stages, integrated within a hierarchical RAN architecture. To validate the proposed framework, a case study demonstrating the utility of the proposed framework using decision Mamba has been presented for the first time in the literature.
@@ II GenAI-based Intent-driven Network Management Schemes
Intent-based networking connects end-user service requests with network operations. Nijah et al. states that NLP-based methods offer the most flexibility to enable intent expression in natural language [2]. LLMs have revolutionized NLP and they surpass traditional methods in flexibility and accuracy. Their ability to perform zero-shot, few-shot, and fine-tuned learning makes them state-of-the-art for tasks like intent recognition and processing. Existing research on IDN management schemes vastly uses LLMs to understand intents from the operators. While a few GenAI paradigms explore techniques like diffusion models and neuro-symbolic reasoning, the majority of current research in IDN management focuses on LLMs. Therefore, in this section, we review the LLM-centric approaches for IDN management in the telecommunication domain. Subsequently, we explore the potential application of LLM techniques or GenAI algorithms which to date, have not been widely explored in the literature. We also summarize the LLM-centric IDN management approaches, their main features, challenges, and applications in Table I.
@@
LLM-assisted intent processing and management for Fifth-Generation (5G) core networks [9, 10] Supports multiple intent transmission to the model. Enhances network orchestration with semantic routing. Prevents LLM hallucination. Includes a RAG module. Reliance on LLMs which are not inherently designed for network control and automation. Distinguishing between similar or overlapping intents. Conversion of operator-defined intents into slice deployment policies. Mapping intents to edge-based service deployments.
LLM-Driven multi-agent and negotiation-based intent management for 6G networks [11, 5] Employs MARL and LLM-driven negotiation. Resolves conflicting intents in multi-stakeholder settings. Cross-domain conflict resolution. Fairness in resource allocation. Scalability concerns. 6G network slicing. Dynamic spectrum allocation. Autonomous resource management.
End-to-end AI-based intent-driven network automation [1, 2, 4] Converts high-level intents into actionable network policies. Verifies if an intent is feasible given the network’s current or predicted state. End-to-end automation using AI. Computational complexity of processing and validating multiple intents in real-time increases exponentially. Zero-touch network configuration. Predictive maintenance and fault recovery.
Generative AI and multimodal intent-based network management [12] Uses multimodal generative AI for intent translation. LLMs are fine-tuned with few-shot learning to adapt to networking use cases. High dependency on predefined templates. Expansion to include multimodal inputs (e.g., images, topology files). Automation of the deployment and management of network slices.
II-A State-of-the-art LLM-Based Approaches for IDN Management
@@ II-A1 LLM-Assisted Intent Processing and Management for 5G Core Networks
Two recent studies explore LLM-assisted approaches on intent-based management in 5G core networks. Manias et al. in their work [9] explore LLM-driven intent extraction as a key enabler of next-generation zero-touch network service management. A customized LLM model is designed to interpret and translate user intents into actionable network policies. By utilizing LLMs, the framework reduces human intervention while maintaining adaptability across diverse network management tasks, such as deployment, modification, performance assurance, and feasibility checking.
Semantic routing is introduced to refine LLM-assisted intent-based networking in [10]. Traditional LLM-driven approaches face issues such as hallucination, scalability limitations, and accuracy degradation when handling complex network intents. To address these challenges, the proposed framework integrates a semantic router, which ensures deterministic decision-making by routing extracted intents through predefined pathways rather than relying solely on LLM-generated responses. This hybrid architecture improves reliability, reduces latency, and enhances accuracy in 5G core management and orchestration.
@@ II-A2 LLM-Driven Multi-Agent and Negotiation-Based Intent Management
Another study [5] introduces a comprehensive LLM-driven intent life cycle management system designed to handle all stages of intent processing, from decomposition and translation to negotiation, activation, and assurance. By shifting away from rigid JSON/YAML intent definitions, the system proposed in [5] enables natural language-driven network orchestration. Users can specify operational goals without requiring deep technical expertise.
@@ II-A3 End-to-End AI-Based Intent-Driven Network Automation
Recent studies introduce end-to-end AI frameworks that integrate LLMs, algorithms like Deep Reinforcement Learning (DRL), Multi-Agent RL (MARL), and Hierarchical RL (HRL) to realize fully automated network management.
While processing intents is crucial, ensuring their feasibility and impact is equally important [4]. A recent study introduces a transformer-based time series predictor to validate intents before execution. This predictive validation module analyzes historical network data and forecasts traffic patterns to ensure that requested optimizations (e.g., increasing energy efficiency, improving throughput) do not negatively affect service quality. Once an intent is validated, an HRL framework selects and triggers appropriate network optimization applications such as beamforming, traffic steering, and power control. An attention-based HRL model filters out suboptimal actions to reduce computational overhead while maximizing efficiency.
Another breakthrough in end-to-end AI-driven network automation is the integration of multi-agent learning frameworks [1], where AI-driven network agents negotiate and resolve conflicting intents in real-time. This is particularly essential in multi-tenant 6G networks, where various stakeholders (e.g., operators, service providers, enterprise users) have competing resource demands. The study conducted in [1] proposes a multi-domain orchestration model, where LLM-driven agents representing RAN, core, and transport domains collaborate to optimize resource allocation and maintain Quality-of-Service (QoS).
@@ II-A4 Generative AI and Multimodal Intent-Based Network Management
A multimodal intent recognition framework presented in [12] utilizes LLMs with GenAI to process diverse inputs like text, images, and deployment descriptors to enable network operators to specify intents through text prompts, graphical designs, or configuration files. AI dynamically refines intents to ensure accuracy while aligning with service-level objectives.
@@
II-B Potential GenAI Paradigms for Enhancing IDN Management
The key limitations of the LLM-based IDN management schemes that we discussed in this section can be described as follows: LLM hallucination and interpretability challenges, which can result in inaccurate or overly generalized outputs; absence of a real-time feedback mechanism, making it difficult to correct errors in network policies after deployment; and deployment challenges, particularly due to memory constraints and high computational requirements. To address these issues, we recommend some possible solutions from the GenAI perspective.
$\bullet$ Retrieval Augmented Generation (RAG) and separate intent refinement/ validation scheme: We can use network-specific databases and retrieval models to ensure LLMs only generate responses based on verified knowledge to reduce hallucination. Furthermore, a multi-step intent refinement process can be introduced where a second model (e.g., a smaller LLM or rule-based checker) cross-checks intent translations before execution.
$\bullet$ Reinforcement Learning with Human Feedback (RLHF): Since re-training is not always an option due to computational resource scarcity, RLHF can be a powerful technique for improving LLM-based IDN management to address the issue of real-time adaptation and feedback loops. We can define a reward function based on network Key Performance Indicators (KPIs) (e.g., latency, throughput, energy efficiency). Then, a feedback loop can be implemented where operators and simulated environments score the LLM ’s intent outputs. Lastly, RL algorithms can be used to refine LLM-generated responses.
@@ II-B Potential GenAI Paradigms for Enhancing IDN Management
III Three-step Methodology for Intent-driven Network Management
In this section of the paper, we present a three-step framework to introduce IDN management in modern-day communication systems. In particular, we propose a three-fold strategy consisting of intent processing, intent validation, and intent execution via network optimization policy generation.

### v2

# Generative AI for Intent-Driven Network Management in 6G RAN: A Case Study on the Mamba Model

@@ Abstract
With the emergence of 6G, mobile networks are becoming increasingly heterogeneous and dynamic, necessitating advanced automation for efficient management. Intent-Driven Networks (IDNs) address this by translating high-level intents into optimization policies. Large Language Models (LLMs) can enhance this process by understanding complex human instructions, enabling adaptive and intelligent automation. Given the rapid advancements in Generative AI (GenAI), a comprehensive survey of LLM-based IDN architectures in disaggregated Radio Access Network (RAN) environments is both timely and critical. This article provides such a survey, along with a case study on a selective State-Space Model (SSM)-enabled IDN architecture that integrates GenAI across three key stages: intent processing, intent validation, and intent execution. For the first time in the literature, we propose a hierarchical framework built on Mamba-SSM that introduces GenAI across all stages of the IDN pipeline. We further present a case study demonstrating that the proposed Mamba architecture significantly improves network performance through intelligent automation, surpassing existing IDN approaches. In a multi-cell 5G/6G scenario, the proposed architecture reduces quality of service drift by up to $70\%$, improves throughput by up to $80$ Mbps, and lowers inference time to 60607070 ms, outperforming GenAI, reinforcement learning, and non-machine learning baselines.
@@ I Introduction
Sixth-Generation (6G) networks are anticipated to support a diverse set of user requirements and have more complex deployments [1]. Traditional network management methods, which are based on manual configurations and human expertise, will face challenges when optimizing and operating such complex networks. Furthermore, manual configurations are costly, prone to errors and not scalable [1]. Intent-Driven Network (IDN) management is emerging as a solution to the challenges [1].
An intent defines a desired network outcome by specifying target metrics without detailing the underlying implementation steps, thereby ensuring both human comprehensibility and machine interpretability [3]. Advances in Natural Language Processing (NLP), driven primarily by the rise of Large Language Models (LLMs) have enabled operators to express intents more naturally and effectively. Beyond intent processing via LLMs, recent works have incorporated other forms of Generative Artificial Intelligence (GenAI) across the IDN pipeline, including transformer-based traffic prediction for intent validation [3] and policy execution via Decision Transformers [19]. These GenAI-enabled components in IDN management collectively reduce manual workload, lower configuration-related errors, and accelerate service deployment and adaptation [4].
However, an important limitation in existing works is the adoption of transformer-based GenAI models at every stage of IDN, including intent processing, predictive validation, and policy execution. This introduces a quadratic complexity bottleneck that increases inference latency and limits scalability as sequence lengths grow. To address this issue, a selective State Space Model (SSM) named Mamba is selected. Mamba provides a suitable alternative by enabling linear-time inference and efficient long-range dependency modeling.
Mamba extends the SSM family by maintaining a recurrent hidden state while replacing fixed linear dynamics with a selective state-update mechanism. This design replaces the Transformer’s quadratic self-attention with a fast and selectively updated memory stream [9]. In LLMs, this enables long-context reasoning without computational explosion and supports efficient processing of large text inputs, logs, and multi-page instructions. In decision-making settings, such as Decision Mamba for Reinforcement Learning (RL) [16], selective memory retains only historically meaningful transitions and avoids uniform weighting of past states. In time-series forecasting, Mamba captures both smooth temporal trends and sudden spikes without relying on positional encodings that limit extrapolation [20]. Across these domains (LLM, time-series forecasting, and control decision making), Mamba addresses the structural limitations of Transformers while offering a faster, lighter, and more stable alternative.
Motivated by the need for more efficient GenAI techniques in hierarchical and disaggregated Radio Access Network (RAN) environments, this work introduces Mamba as an alternative to transformer-based algorithms for IDN management. We begin with a brief review of state-of-the-art GenAI methods used across the IDN workflow. We then propose a three-stage Mamba-driven methodology that covers intent processing, validation, and execution. Building on this methodology, we also show how each stage can be realized within a hierarchical RAN architecture through multiple specialized Mamba-based models. A case study further demonstrates the practical realization of the entire pipeline. Specifically, intents are interpreted using a parameter-efficient Mamba-based LLM fine-tuned on a custom dataset [21]. Intent feasibility is assessed using a Mamba-based time-series predictor that forecasts key network parameters and validates operator objectives accordingly. Finally, intent execution is performed via Decision Mamba, for its superior memory and computational efficiency compared to transformer-based approaches [9].
The core contributions of this paper are threefold. First, it presents a precise, yet comprehensive survey of GenAI-based IDN architectures, highlighting their main features, limitations and applications. Along with the survey, we also present open research directions, practical deployment considerations, and standardization implications for GenAI-enabled IDN management, offering a roadmap for future development in this domain. Second, we introduce the first Mamba-driven IDN framework in the literature and demonstrate how it can replace traditional transformer-based components to deliver linear-time inference with improved long-range reasoning. Third, we show the effectiveness of the proposed framework through a detailed case study within a hierarchical RAN architecture that showcases the end-to-end workflow of Mamba-based IDN management and highlights its computational and operational advantages.
Unlike prior GenAI-enabled IDN frameworks that rely on transformer-based models and typically introduce intelligence at only one or two stages of the IDN pipeline, this work is the first to systematically replace transformers across intent processing, validation, and execution using the Mamba models. This unified Mamba-driven architecture eliminates quadratic attention overhead, enables faster inference, and improves crucial Key Performance Indicators (KPIs).
In our evaluation across a multi-cell Fifth-Generation (5G)/6G scenario, the proposed Mamba–enabled IDN architecture delivers substantial performance gains across all major KPIs. The framework consistently achieves the lowest Quality-of-Service (QoS) drift, reducing intent-violation rates by up to $70\%$ compared to the baselines. It further enhances network throughput by 45458080 Mbps under moderate to high traffic loads and achieves up to two times higher energy efficiency relative to the baselines. In addition, the proposed Mamba architecture for IDN management enables fast inference, lowering decision latency to 60607070 ms, which is considerably better than the baselines as well.
@@ II GenAI-Based Intent-driven Network Management Schemes
LLMs have transformed IDN management due to their ability to perform zero-shot, few-shot, and fine-tuned learning. They now represent the state-of-the-art for intent recognition and processing, and most existing IDN studies employ them to interpret operator intents. The literature has also evolved beyond intent processing, incorporating AI-based modules for validating and executing intents [1, 3]. Recently, a full-scale GenAI-driven solution was proposed in [19]. Motivated by these developments, this section first reviews works that focus specifically on GenAI-based methods directly supporting the IDN management workflow across intent processing, validation, and execution. We summarize reviewed works in terms of their main features, challenges, and applications in Table I. Note that we have selected works that have used GenAI in at least one of the stages of the IDN management workflow. Subsequently, in this section, we explore the open research directions, 3GPP and Open RAN standard implications, and practical deployment considerations for GenAI-based IDN management.
@@
LLM-assisted intent processing and management for 5G core networks [11, 12] Supports multiple intent transmission to the model. Enhances network orchestration with semantic routing. Prevents LLM hallucination. Includes a Retrieval Augmented Generation (RAG) module. Reliance on LLMs, which are not inherently designed for network control and automation. Distinguishing between similar or overlapping intents. Conversion of operator-defined intents into slice deployment policies. Mapping intents to edge-based service deployments.
LLM-Driven multi-agent and negotiation-based intent management for 6G networks [13, 4] Employs Multi-Agent RL (MARL) and LLM-driven negotiation. Resolves conflicting intents in multi-stakeholder settings. Cross-domain conflict resolution. Fairness in resource allocation. Scalability concerns. 6G network slicing. Dynamic spectrum allocation. Autonomous resource management.
End-to-end AI-based intent-driven network automation [1, 3, 19, 22] Converts high-level intents into actionable network policies. Verifies if an intent is feasible given the network’s current or predicted state. End-to-end automation using AI. Computational complexity of processing and validating multiple intents in real-time increases exponentially. Zero-touch network configuration. Predictive maintenance and fault recovery.
Generative AI and multimodal intent-based network management [15] Uses multimodal generative AI for intent translation. LLMs are tuned with few-shot learning. High dependency on predefined templates. Expansion to include multimodal inputs. Automation of the deployment and management of network slices.
II-A State-of-the-art GenAI-Based Approaches for IDN Management
@@ II-A1 LLM-Assisted Intent Processing and Management for 5G Core Networks
Two recent studies explore LLM-assisted approaches on intent-based management in 5G core networks. Manias et al. in their work [11] explore LLM-driven intent extraction as a key enabler of next-generation zero-touch network service management. A customized LLM model is designed to interpret and translate user intents into actionable network policies. By utilizing LLMs, the framework reduces human intervention while maintaining adaptability across diverse network management tasks.
Semantic routing is introduced to refine LLM-assisted intent-based networking in [12]. Traditional LLM-driven approaches face issues such as hallucination, scalability limitations, and accuracy degradation when handling complex network intents. To address these challenges, the proposed framework integrates a semantic router, which ensures deterministic decision-making by routing extracted intents through predefined pathways rather than relying solely on LLM-generated responses.
@@ II-A2 LLM-Driven Multi-Agent and Negotiation-Based Intent Management
Another study [4] introduces a comprehensive LLM-driven intent life cycle management system designed to handle all stages of intent processing, from decomposition and translation to negotiation, activation, and assurance.
@@ II-A3 End-to-End AI-Based Intent-Driven Network Automation
Recent studies introduce end-to-end AI frameworks that integrate LLMs, algorithms like Deep Reinforcement Learning (DRL), MARL, and Hierarchical RL (HRL) to realize fully automated network management.
A recent study introduces a transformer-based time series predictor to validate intents before execution [3]. This predictive validation module analyzes historical network data and forecasts traffic patterns to ensure that requested optimizations do not negatively affect service quality. Once an intent is validated, an HRL framework selects and triggers appropriate network optimization applications such as beamforming, traffic steering, and power control. An attention-based HRL model filters out suboptimal actions to reduce computational overhead while maximizing efficiency.
Zhang et al. propose an intent-driven closed-loop control and management framework for 6G Open RAN, which integrates AI across all stages of the IDN pipeline [22]. Intent translation is performed using a long short-term memory model with morphological rules, while intent verification and decomposition rely on AI-assisted reasoning via event calculus. The refined objectives feed into a control loop, where a Deep-Q-Network (DQN)-based network application (near-real-time control) executes slice-level resource allocation.
Another recent contribution is proposed in [19], which demonstrates a fully GenAI-enabled workflow for intent processing, validation, and execution. The framework integrates a Quantized Low Rank Adaptation (QLoRA)-fine-tuned LLM for intent interpretation, an Informer model for long-horizon traffic prediction, and a goal-aware Decision Transformer (DT) for RAN control. By combining large-scale language understanding with predictive analytics and sequential decision-making, this work illustrates how GenAI can support seamless end-to-end automation in 6G Open RAN environments.
Another breakthrough in end-to-end AI-driven network automation is the integration of multi-agent learning frameworks [1], where AI-driven network agents negotiate and resolve conflicting intents in real-time. This is particularly essential in multi-tenant 6G networks, where various stakeholders have competing resource demands.
@@ II-A4 Generative AI and Multimodal Intent-Based Network Management
A multimodal intent recognition framework presented in [15] utilizes LLMs with GenAI to process diverse inputs like text, images, and deployment descriptors to enable network operators for specifying intents through text prompts, graphical designs, or configuration files. AI dynamically refines intents to ensure accuracy while aligning with service-level objectives.
@@
II-B Open Research Directions for Enhancing IDN Management
The key limitations of the GenAI-based IDN management schemes discussed so far in this section can be described as follows: LLM hallucination and interpretability challenges, which can result in inaccurate or overly generalized outputs; absence of a real-time feedback mechanism, making it difficult to correct errors in network policies after deployment; and deployment challenges, particularly due to memory constraints and high computational requirements. To address these issues, we recommend some possible measures as open research directions.
$\bullet$ RAG: We can use network-specific databases and retrieval models to ensure LLMs only generate responses based on verified knowledge to reduce hallucination. Particularly, there has not been any work in the IDN literature that has explored Agentic RAG (using LLM agent to choose between databases).
$\bullet$ Reinforcement Learning with Human Feedback (RLHF): Since re-training is not always an option due to computational resource scarcity, RLHF can be a powerful technique for improving LLM-based IDN management to address the issue of real-time adaptation and feedback loops. We can define a reward function based on network KPIs (e.g., latency, throughput, energy efficiency). Then, a feedback loop can be implemented where operators and simulated environments score the LLM ’s intent outputs. Lastly, RL algorithms can be used to refine LLM-generated responses.
@@ II-B Open Research Directions for Enhancing IDN Management
$\bullet$ Mamba models: Mamba models can be incorporated for reduced memory overhead and faster inference instead of the transformers in the IDN management pipeline.
II-C Practical deployment considerations
Practical deployment of GenAI-enabled IDN management must account for the strict latency requirements of RAN environments, where many control-loop decisions must be completed within the near-real-time window of 1010ms to $1$s. These constraints limit the size and complexity of deployable LLMs and predictive models. Another key challenge is ensuring robustness against QoS and intent drift, as real-world network conditions and operator objectives may shift over time. To maintain reliable operation, deployment pipelines must incorporate lightweight model updates, drift detection mechanisms, and safeguards that ensure GenAI-driven decisions remain compliant with QoS guarantees.
II-D 3GPP and Open RAN standard implications
GenAI-enabled IDN management has several implications for ongoing 3GPP and Open RAN standardization efforts. First, 3GPP Release 17 and beyond define intent-driven management services (TS 28.312) , but current specifications assume rule-based intent translation. Integrating LLM-based intent processing and SSM-driven validation requires extending these standards to support natural-language intent formats, semantic models, and AI-assisted intent interpretation pipelines. Second, predictive validation and goal-aware decision-making must align with 3GPP’s management data models, where AI-generated recommendations map to standardized network resource models and service exposure Application Programming Interfaces (APIs).
From the Open RAN perspective, introducing GenAI into the intent execution loop affects the interfaces between the service management orchestrator, non-real-time intelligent controller, near-real-time intelligent controller, and network applications inhabiting in those controllers. Mamba-based predictors and decision models must comply with O-RAN open interfaces such as A1 (connects the non and near-real-time controllers), E2 (links the near-real-time controller with the distributed RAN nodes (O-RAN Distributed Unit (O-DU)/O-RAN Centralized Unit (O-CU)) for cross-layer orchestration. Moreover, near-real-time latency requirements (1010ms–11s) impose constraints on the deployment of GenAI models that motivate faster inference and hardware-aware optimization.
III Three-step Methodology for IDN Management and its RAN Integration with GenAI
In this section of the paper, we discuss a three-step framework consisting of intent processing, validation and execution to introduce IDN management and how GenAI can be integrated within it. This section will help the readers have the necessary background to understand the Mamba-based case study to be presented in the next section.